# Robust Nonlinear Least Squares Problems

### v1

@@ Abstract
This paper considers min-max optimization for a class of robust nonlinear least squares problems. We show via an example that solving the first order optimality conditions defined by gradients of the objective function can lead to incorrect solutions of min-max problems. We give an explicit formula for the value function of the inner maximization problem. Using the formula, we show that finding a first order $\epsilon$-approximate necessary minimax point of the min-max problem needs at most $O(|\log\epsilon|+\epsilon^{-2})$ evaluations of the function value and gradients of the objective function. Moreover, we establish error bounds from any solution of the nonlinear least squares problem to the solution set of the robust nonlinear least squares problem.
keywords:
Min-max optimization, nonlinear least squares problems, robust learning, nonsmooth nonconvex optimization, complexity analysis
{AMS}
49K35, 90C30, 90C31
@@ 1 Introduction
$\min_{\mathbf{x}}\|F(\mathbf{x})\|^{2}$ (1)
is widely used in data analysis and machine learning. For example, Recurrent Neural Netwoks (RNN)[2, 5] are widely used to model sequential data in many areas such that natural language processing, speech recognition, machine translation, and stock market analysis. Given input and output data sets $\{\mathbf{a}_{i}\in\mathbb{R}^{\hat{n}}\}_{i=1}^{N}$ and $\{\mathbf{b}_{i}\in\mathbb{R}^{k}\}_{i=1}^{N}$, the training process of RNN is represented as
$\min_{\mathbf{x}}\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{b}_{i}-\mathbf{U}\sigma(\mathbf{W}(\cdots\sigma(\mathbf{V}\mathbf{a}_{1}+\mathbf{u})\cdots)+\mathbf{V}\mathbf{a}_{i}+\mathbf{u})+\mathbf{v})\|^{2},$ (2)
where $\mathbf{U}\in\mathbb{R}^{k\times l}$, ${\bf W}\in\mathbb{R}^{l\times l}$, ${\bf V}\in\mathbb{R}^{l\times\hat{n}}$ are the weight matrices, $\mathbf{u}\in\mathbb{R}^{l},\mathbf{v}\in\mathbb{R}^{k}$ are the bias vectors, $\sigma$ is an activation function, e.g., the ReLU activation function $\sigma({\bf z})=\max(0,{\bf z})$, and
$\mathbf{x}=({\rm vec}(\mathbf{U})^{\top},{\rm vec}(\mathbf{V})^{\top},{\rm vec}(\mathbf{W})^{\top},\mathbf{u}^{\top},\mathbf{v}^{\top})^{\top}\in\mathbb{R}^{n}$
with $n={kl+l^{2}+l\hat{n}+l+k}$ and $m=Nk$. When we study a big set of data, $m=Nk$ can be much bigger than $n$.
Problem (2) is an overdetermined least squares problem. Another example of overdetermined least squares problems is training convolutional neural network(CNN). For example, a two-layer CNN [3, 13]
$\min_{\mathbf{x}}\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{b}_{i}-\mathbf{U}\sigma(\mathbf{W}\sigma(\mathbf{V}\mathbf{a}_{i}+\mathbf{u})+\mathbf{v})\|^{2}.$ (3)
When the output data $\{\mathbf{b}_{i}\in\mathbb{R}^{k}\}_{i=1}^{N}$ contain noise, the following min-max problem
@@
is used for robust model training [17], where $\mathbf{C}\in\mathbb{R}^{m\times r}$ with $r\leq m$, $\Omega=\{\mathbf{y}\,:\,\|\mathbf{y}\|_{\infty}\leq\delta\}$, and $\delta\geq 0$. If $\delta=0$, then problem (4) reduces to problem (1).
First order methods including decent-ascent methods have been widely used to find the first order stationary points of min-max problems [4, 11, 10, 17]. However, the first order stationary points defined by gradients of the objective function $f$ can lead to incorrect solutions of min-max problems. In subsection 1.2, we show with an example of robust linear least square problems that solving the first order optimality conditions can lead to incorrect results.
@@
    for any $\mathbf{x}\in\mathbb{R}^{n}$, where
    $\|{\bf C}\|_{F}=\left(\sum_{i=1,j=1}^{m,r}{\bf C}_{ij}^{2}\right)^{1/2}$
    is the Frobenius norm of ${\bf C}$.
  • We apply the formula for $\varphi$ to show that the method proposed by Gratton, Simon and Toint [9] can be used to find a first order $\epsilon$-approximate necessary minimax point of problem (4) with optimal complexity.
  • We give error bounds from any solution of nonlinear least-squares problem (1) to the set of minimizers of $\varphi.$
@@ 1.1 Important Concepts
A point $(\mathbf{x}^{*},\mathbf{y}^{*})\in\mathbb{R}^{n}\times\Omega$ is called a saddle point [16] of problem (4) if for any $\mathbf{x}\in\mathbb{R}^{n},\mathbf{y}\in\Omega,$ we have
@@
A point $(\mathbf{x}^{*},\mathbf{y}^{*})\in\mathbb{R}^{n}\times\Omega$ is called a minimax point [11] of problem (4) if
@@
By Theorem 3.11 in [10], if $f$ is directionally differentiable then a local minimax point $(\mathbf{x}^{*},\mathbf{y}^{*})\in\mathbb{R}^{n}\times\Omega$ of problem (4) satisfies the following first order condition
$\begin{array}[]{ll}d_{\mathbf{x}}f(\mathbf{x}^{*},\mathbf{y}^{*};\mathbf{v})=0,\quad\forall\,\mathbf{v}\in\mathbb{R}^{n}\\ d_{\mathbf{y}}f(\mathbf{x}^{*},\mathbf{y}^{*};\mathbf{u})\leq 0,\quad\forall\,\mathbf{u}\in\mathrm{T}_{\Omega}(\mathbf{y}^{*}),\end{array}$ (7)
where $d_{\mathbf{x}}(\mathbf{x}^{*},\mathbf{y}^{*};\mathbf{v})$ and $d_{\mathbf{y}}(\mathbf{x}^{*},\mathbf{y}^{*};\mathbf{u})$ denote the directional derivatives of $f$ at $(\mathbf{x}^{*},\mathbf{y}^{*})$ along the directions $\mathbf{v}$ and $\mathbf{u}$, respectively, and $\mathrm{T}_{\Omega}(\mathbf{y})=:\left\{\mathbf{w}:\exists~{}\mathbf{y}^{k}\overset{\Omega}{\to}\mathbf{y},t^{k}\downarrow 0~{}\text{such that}~{}\lim_{k\to\infty}\frac{\mathbf{y}^{k}-\mathbf{y}}{t^{k}}=\mathbf{w}\right\}$ is the tangent cone to $\Omega$ at $\mathbf{y}\in\Omega$ [15, Definition 6.1]. When $\mathbf{F}$ is differentiable, the first order optimality condition of problem (4) has the following form
@@
is nonempty and bounded for any $\mathbf{x}\in\mathbb{R}^{n}$. Hence, the function $\varphi$ is well-defined. However, it is known that solving a convex quadratic maximization over a box feasible set is NP-hard [14]. People often consider the first order optimality condition for nonconvex-nonconcave min-max problem [4, 10, 17]. However a solution of this first order optimality condition can be meaningless for (4).
@@ 1.2 Example
Consider problem (4) with $\mathbf{F}(\mathbf{x})=\mathbf{A}\mathbf{x}-\mathbf{b}$, $\mathbf{C}=\mathbf{I}$ and $n=m=r$. Suppose that $\mathbf{A}\in\mathbb{R}^{n\times n}$ is nonsingular. Then any pair $(\mathbf{x},\mathbf{y})=(\mathbf{A}^{-1}(\mathbf{b}+\mathbf{y}),\mathbf{y})$ with $\mathbf{y}\in\Omega$ satisfies the first order condition (10). On the other hand, the optimal solution of the inner maximization problem has the closed form
$\mathbf{y}_{i}(\mathbf{x})=\mathop{\mathrm{argmax}}_{|\mathbf{y}_{i}|\leq\delta}(\mathbf{A}\mathbf{x}-\mathbf{b}-\mathbf{y})_{i}^{2}=\delta\left\{\begin{array}[]{ll}1&{\rm if}\,(\mathbf{A}\mathbf{x}-\mathbf{b})_{i}<0\\ \{-1,1\}&{\rm if}\,(\mathbf{A}\mathbf{x}-\mathbf{b})_{i}=0\\ -1&{\rm if}\,(\mathbf{A}\mathbf{x}-\mathbf{b})_{i}>0,\end{array}\right.\quad i=1,\ldots,r.$ (11)
@@
$\min_{\mathbf{x}}\max_{\mathbf{y}\in\Omega}\|\mathbf{A}\mathbf{x}-\mathbf{b}-\mathbf{y}\|^{2}=\min_{\mathbf{x}}\|\mathbf{A}\mathbf{x}-\mathbf{b}-\mathbf{y}(\mathbf{x})\|^{2}=r\delta^{2}$
at $(\mathbf{x}^{*},\hat{\mathbf{y}})=(\mathbf{A}^{-1}\mathbf{b},\delta\mathbf{c})$, where $\mathbf{c}$ is a vector with all elements being 1 or -1. In fact, by the Danskin theorem, the inner function $\|\mathbf{A}\mathbf{x}-\mathbf{b}-\mathbf{y}(\mathbf{x})\|^{2}$ is strongly convex. Hence, the minimization problem has only one solution $\mathbf{x}^{*}=\mathbf{A}^{-1}\mathbf{b}.$ It is worth noting that $(\mathbf{x}^{*},\hat{\mathbf{y}})$ is not a saddle point of $f(\mathbf{x},\mathbf{y})=\|\mathbf{A}\mathbf{x}-\mathbf{b}-\mathbf{y}\|^{2},$ for any $\hat{\mathbf{y}}$ with $\hat{\mathbf{y}}_{i}\in\{\delta,-\delta\}$, since
$f(\mathbf{x}^{*},\hat{\mathbf{y}})=r\delta^{2}>f(\mathbf{A}^{-1}(\mathbf{b}+\hat{\mathbf{y}}),\hat{\mathbf{y}})=0.$
@@
$r\delta^{2}=\min_{\mathbf{x}}\max_{\mathbf{y}\in\Omega}f(\mathbf{x},\mathbf{y})\neq\max_{\mathbf{y}\in\Omega}\min_{\mathbf{x}}f(\mathbf{x},\mathbf{y})=0.$
See [6, Theorem 1.4.1]. Moreover, due to the arbitrary choice of $\delta$ for the closed form solution (11) of the maximization problem, we can easily find that $(\mathbf{x}^{*},\hat{\mathbf{y}})$ is not a local saddle point.
@@
Hence $(\mathbf{x}^{*},\hat{\mathbf{y}})$ is a minimax point of $f(\mathbf{x},\mathbf{y})=\|\mathbf{A}\mathbf{x}-\mathbf{b}-\mathbf{y}\|^{2}$. However, $(\mathbf{x}^{*},\hat{\mathbf{y}})$ is not a local minimax point of $f$. Indeed, consider $\hat{\mathbf{y}}=\delta\mathbf{e}$, where $\mathbf{e}$ is the vector with all elements being 1.
@@
$|{\bf A}\mathbf{x}-\mathbf{b}|_{i}\leq\|{\bf A}\mathbf{x}-\mathbf{b}\|=\|{\bf A}(\mathbf{x}-\mathbf{x}^{*})\|\leq\|{\bf A}\|\rho\leq\delta-\frac{\tau(\rho)}{2}$
and for $i=1\ldots,r$,
@@
Let $\mathbf{x}^{1}={\bf A}^{-1}(\mathbf{b}+\epsilon\mathbf{e})$ with $\epsilon\leq\rho/\|{\bf A}^{-1}\mathbf{e}\|$. Then we obtain $\|\mathbf{x}^{1}-\mathbf{x}^{*}\|\leq\rho$ and
$r\delta^{2}=f(\mathbf{x}^{*},\hat{\mathbf{y}})$ $>$ $\max_{\|\mathbf{y}\|_{\infty}\leq\delta,\|\mathbf{y}-\delta\mathbf{e}\|_{\infty}\leq\tau(\rho)}\|\mathbf{A}\mathbf{x}^{1}-\mathbf{b}-\mathbf{y}\|^{2}=\|\epsilon\mathbf{e}-\mathbf{y}(\mathbf{x})\|^{2}$
$=$ $(\epsilon-\delta)^{2}r\geq\min_{\mathbf{x}}\max_{\|\mathbf{y}\|_{\infty}\leq\delta,\|\mathbf{y}-\delta\mathbf{e}\|_{\infty}\leq\tau(\rho)}f(\mathbf{x},\mathbf{y}).$
Hence, $(\mathbf{x}^{*},\hat{\mathbf{y}})$ is not a local minimax point of $f$. Moreover, it is easy to see that $(\mathbf{x}^{*},\hat{\mathbf{y}})$ does not satisfy the first order condition (10), since ${\bf A}^{\top}({\bf A}\mathbf{x}^{*}-{\bf b}-\hat{\mathbf{y}})=-\delta{\bf A}^{\top}\mathbf{e}\neq 0$. On the other hand, if $(\mathbf{x}^{\prime},\mathbf{y}^{\prime})$ satisfies the first order condition (10), then from the nonsingularity of ${\bf A}$, we have $f(\mathbf{x}^{\prime},\mathbf{y}^{\prime})=0$, which implies $(\mathbf{x}^{\prime},\mathbf{y}^{\prime})$ cannot be a local minimax point or a minimax point, since $00=f(\mathbf{x}^{\prime},\mathbf{y}^{\prime})\not\geq\max_{\mathbf{y}\in\Omega}f(\mathbf{x}^{\prime},\mathbf{y})>$ 0.
Note that for nonconvex-nonconcave min-max optimization, a local minimax point must satisfy the first order condition, but a minimax point is not necessarily a stationary point[10].
This example shows that a min-max problem for robust linear least squares problems does not have a saddle point, a local saddle point and a local minimax point, while it has first order stationary points and minimax points. However, the set of first order stationary points and the set of minimax points do not have a common point.
This example also shows that the study of the value function of the inner maximization problem is important. In § 2, we give a proof of closed form expression (5) for the inner value function. Using the inner value function, we prove the existence of a minimax point of the robust nonlinear least squares problem (4). In § 3 we apply that formula to obtain a complexity bound for finding a first order stationary point for a minimax point of (4). Finally, in § 4 we apply the formula to expose some connections between solutions of (4) and those of (1), and give error bounds from any solution of nonlinear least squares problem (1) to the set of minimizers of $\varphi.$

### v2

# X. Chen and C.T. Kelley

@@ Abstract
This paper considers robust solutions to a class of nonlinear least squares problems using min-max optimization approach. We give an explicit formula for the value function of the inner maximization problem and show the existence of global minimax points. We establish error bounds from any solution of the nonlinear least squares problem to the solution set of the robust nonlinear least squares problem. Moreover, we propose a smoothing method for finding a global minimax point of the min-max problem by using the formula and show that finding an $\epsilon$ minimax critical point of the min-max problem needs at most $O(\epsilon^{-2}+\delta^{2}\epsilon^{-3})$ evaluations of the function value and gradients of the objective function, where $\delta$ is the tolerance of the noise. Numerical results of integral equations with uncertain data demonstrate the robustness of solutions of our approach and unstable behaviour of least squares solutions disregarding uncertainties in the data. Min-max optimization; nonlinear least squares problems; numerical solution of integral equations; nonsmooth nonconvex optimization, complexity
@@ 1 Introduction
$\min_{\mathbf{x}}\|\mathbf{F}(\mathbf{x})\|^{2}$ (1)
is widely used to model problems in science and engineering. When the problems contain uncertain or noisy data, it is important to find a robust solution. In this paper, we consider the following min-max problem
@@
for the robust nonlinear least squares problem where $\mathbf{C}\in\mathbb{R}^{m\times r}$ with $r\leq m$, $\Omega=\{\mathbf{y}\in\mathbb{R}^{r}\,:\,\|\mathbf{y}\|_{\infty}\leq\delta\}$, and $\delta\geq 0$. If $\delta=0$, then problem (2) reduces to problem (1). In some applications, the perturbation vector has a known structure and can be present as a finite linear combination of columns of $C$. See Ghaoui & Lebret (1997) and §5.2 in this paper.
A particular example is from the first kind Fredholm nonlinear integral equation with noisy data, which can be written in the following form:
[…]
where $h,p,q$ are continuous functions. The low dimensional projection method is a popular numerical method to find an approximate solution of (3). For example, using
$\tilde{u}({\bf x})=\sum^{n}_{i=1}\mathbf{x}_{i}T_{i-1}(x),$
where $T_{i}$ is the $i$th degree Chebyshev polynomial on $[0,1]$, we obtain a nonlinear least squares problem (1). If $h$ contains uncertain or noisy data, we can use the min-max model (2) to find a robust least squares solution.
Nonlinear least square problems are also often seen in data sciences and machine learning, for example, training convolutional neural network (CNN). Given input and output data sets $\{\mathbf{a}_{i}\in\mathbb{R}^{\hat{n}}\}_{i=1}^{N}$ and $\{{\bf{b}}_{i}\in\mathbb{R}^{k}\}_{i=1}^{N}$, the training process of a two-layer CNN (Cui et al., 2020; Liu et al., 2022, 2023) is represented as
$\min_{\mathbf{x}}\frac{1}{N}\sum_{i=1}^{N}\|{\bf{b}}_{i}-\mathbf{U}\sigma(\mathbf{W}\sigma(\mathbf{V}\mathbf{a}_{i}+{\bf{u}})+{\bf{v}})\|^{2},$ (4)
where $\mathbf{U}\in\mathbb{R}^{k\times l}$, ${\bf W}\in\mathbb{R}^{l\times l}$, ${\bf V}\in\mathbb{R}^{l\times\hat{n}}$ are the weight matrices, ${\bf{u}}\in\mathbb{R}^{l},{\bf{v}}\in\mathbb{R}^{k}$ are the bias vectors, $\sigma$ is an activation function, e.g., the ReLU activation function $\sigma({\bf z})=\max(0,{\bf z})$. Let
$\mathbf{x}=({\rm vec}(\mathbf{U})^{\top},{\rm vec}(\mathbf{V})^{\top},{\rm vec}(\mathbf{W})^{\top},{\bf{u}}^{\top},{\bf{v}}^{\top})^{\top}\in\mathbb{R}^{n}$
with $n={kl+l^{2}+l\hat{n}+l+k}$ and $m=Nk$. When we study a big set of data, $m$ can be much bigger than $n$. Problem (4) can be written as problem (1), which is an overdetermined least squares problem. When the output data $\{{\bf{b}}_{i}\in\mathbb{R}^{k}\}_{i=1}^{N}$ contain noise, the min-max problem (2) can be used to find a robust least squares solution.
First order methods including decent-ascent methods have been widely used to find the first order stationary points of min-max problems (Diakonikolas et al., 2021; Jin et al., 2020; Jiang & Chen, 2023; Yang et al., 2020). However, the first order stationary points defined by gradients of the objective function $f$ can lead to incorrect solutions of min-max problems. In subsection 1.2, we show with an example of robust linear least square problems that solving the first order optimality conditions can lead to incorrect results.
@@
    for any $\mathbf{x}\in\mathbb{R}^{n}$, where $\|{\bf C}\|_{F}$ is the Frobenius norm of ${\bf C}$.
  • We propose a smoothing method for finding a global minimax point of the min-max problem by using the function $\varphi$ and show that finding an $\epsilon$ minimax critical point of the min-max problem (2) needs at most $O(\epsilon^{-2}+\delta^{2}\epsilon^{-3})$ evaluations of the function value and gradients of the objective function.
  • We give error bounds from any solution of nonlinear least squares problem (1) to the set of minimizers of $\varphi.$ We give an explicit form for the difference of worst-case residual errors of a least squares solution and a robust least squares solution, and show the difference increases linearly as the tolerance of the noise increases.
@@ 1.1 Important concepts
A point $(\mathbf{x}^{*},\mathbf{y}^{*})\in\mathbb{R}^{n}\times\Omega$ is called a saddle point (v. Neumann, 1928) of problem (2) if for any $\mathbf{x}\in\mathbb{R}^{m},\mathbf{y}\in\Omega,$ we have
@@
A point $(\mathbf{x}^{*},\mathbf{y}^{*})\in\mathbb{R}^{n}\times\Omega$ is called a global minimax point (Jin et al., 2020) of problem (2) if
@@
By Theorem 3.11 in Jiang & Chen (2023), if $f$ is directionally differentiable then a local minimax point $(\mathbf{x}^{*},\mathbf{y}^{*})\in\mathbb{R}^{n}\times\Omega$ of problem (2) satisfies the following first order condition
$\begin{array}[]{ll}d_{\mathbf{x}}f(\mathbf{x}^{*},\mathbf{y}^{*};{\bf{v}})=0,\quad\forall\,{\bf{v}}\in\mathbb{R}^{n}\\ d_{\mathbf{y}}f(\mathbf{x}^{*},\mathbf{y}^{*};{\bf{u}})\leq 0,\quad\forall\,{\bf{u}}\in\mathrm{T}_{\Omega}(\mathbf{y}^{*}),\end{array}$ (7)
where $d_{\mathbf{x}}f(\mathbf{x}^{*},\mathbf{y}^{*};{\bf{v}})$ and $d_{\mathbf{y}}f(\mathbf{x}^{*},\mathbf{y}^{*};{\bf{u}})$ denote the directional derivatives of $f$ at $(\mathbf{x}^{*},\mathbf{y}^{*})$ along the directions ${\bf{v}}$ and ${\bf{u}}$, respectively, and
$\mathrm{T}_{\Omega}(\mathbf{y})=:\left\{\mathbf{w}:\exists~{}\mathbf{y}^{k}\overset{\Omega}{\to}\mathbf{y},t^{k}\downarrow 0~{}\text{such that}~{}\lim_{k\to\infty}\frac{\mathbf{y}^{k}-\mathbf{y}}{t^{k}}=\mathbf{w}\right\}$
is the tangent cone to $\Omega$ at $\mathbf{y}\in\Omega$ (Rockafellar & Wets, 2009, Definition 6.1). When $\mathbf{F}$ is differentiable, the first order optimality condition of problem (2) has the following form
@@
is nonempty and bounded for any $\mathbf{x}\in\mathbb{R}^{n}$. Hence, the function $\varphi$ is well-defined. However, it is known that solving a convex quadratic maximization over a box feasible set is NP-hard (Pardalos & Vavasis, 1991). People often consider the first order optimality condition for nonconvex-nonconcave min-max problem (Diakonikolas et al., 2021; Jiang & Chen, 2023; Yang et al., 2020). However a solution of this first order optimality condition can be meaningless for (2), unless the problem has a local minimax problem, for example, convex-concave saddle point problems with cardinality penalties Bian & Chen (2024).
@@ 1.2 Example
Consider problem (2) with $\mathbf{F}(\mathbf{x})=\mathbf{A}\mathbf{x}-{\bf{b}}$ and $\mathbf{C}=\mathbf{I}$. Suppose that $\mathbf{A}\in\mathbb{R}^{n\times n}$ is nonsingular. Then any pair $(\mathbf{x},\mathbf{y})=(\mathbf{A}^{-1}({\bf{b}}+\mathbf{y}),\mathbf{y})$ with $\mathbf{y}\in\Omega$ satisfies the first order condition (10). On the other hand, the optimal solution of the inner maximization problem has the closed form
$\mathbf{y}_{i}(\mathbf{x})=\mathop{\mathrm{argmax}}_{|\mathbf{y}_{i}|\leq\delta}(\mathbf{A}\mathbf{x}-{\bf{b}}-\mathbf{y})_{i}^{2}=\delta\left\{\begin{array}[]{ll}1&{\rm if}\,(\mathbf{A}\mathbf{x}-{\bf{b}})_{i}<0\\ \{-1,1\}&{\rm if}\,(\mathbf{A}\mathbf{x}-{\bf{b}})_{i}=0\\ -1&{\rm if}\,(\mathbf{A}\mathbf{x}-{\bf{b}})_{i}>0,\end{array}\right.\quad i=1,\ldots,n.$ (11)
@@
$\min_{\mathbf{x}}\max_{\mathbf{y}\in\Omega}\|\mathbf{A}\mathbf{x}-{\bf{b}}-\mathbf{y}\|^{2}=\min_{\mathbf{x}}\|\mathbf{A}\mathbf{x}-{\bf{b}}-\mathbf{y}(\mathbf{x})\|^{2}=n\delta^{2}$
at $(\mathbf{x}^{*},\hat{\mathbf{y}})=(\mathbf{A}^{-1}{\bf{b}},\delta{\bf{c}})$, where ${\bf{c}}$ is a vector with all elements being 1 or -1. In fact, by the Danskin theorem, the inner function $\|\mathbf{A}\mathbf{x}-{\bf{b}}-\mathbf{y}(\mathbf{x})\|^{2}$ is strongly convex. Hence, the minimization problem has only one solution $\mathbf{x}^{*}=\mathbf{A}^{-1}{\bf{b}}.$ It is worth noting that $(\mathbf{x}^{*},\hat{\mathbf{y}})$ is not a saddle point of $f(\mathbf{x},\mathbf{y})=\|\mathbf{A}\mathbf{x}-{\bf{b}}-\mathbf{y}\|^{2},$ for any $\hat{\mathbf{y}}$ with $\hat{\mathbf{y}}_{i}\in\{\delta,-\delta\}$, since
$f(\mathbf{x}^{*},\hat{\mathbf{y}})=n\delta^{2}>f(\mathbf{A}^{-1}({\bf{b}}+\hat{\mathbf{y}}),\hat{\mathbf{y}})=0.$
@@
$n\delta^{2}=\min_{\mathbf{x}}\max_{\mathbf{y}\in\Omega}f(\mathbf{x},\mathbf{y})\neq\max_{\mathbf{y}\in\Omega}\min_{\mathbf{x}}f(\mathbf{x},\mathbf{y})=0.$
See Facchinei & Pang (2002, Theorem 1.4.1). Moreover, due to the arbitrary choice of $\delta$ for the closed form solution (11) of the maximization problem, we can easily find that $(\mathbf{x}^{*},\hat{\mathbf{y}})$ is not a local saddle point.
@@
Hence $(\mathbf{x}^{*},\hat{\mathbf{y}})$ is a global minimax point of $f(\mathbf{x},\mathbf{y})=\|\mathbf{A}\mathbf{x}-{\bf{b}}-\mathbf{y}\|^{2}$. However, $(\mathbf{x}^{*},\hat{\mathbf{y}})$ is not a local minimax point of $f$. Indeed, consider $\hat{\mathbf{y}}=\delta\mathbf{e}$, where $\mathbf{e}$ is the vector with all elements being 1.
@@
$|{\bf A}\mathbf{x}-{\bf{b}}|_{i}\leq\|{\bf A}\mathbf{x}-{\bf{b}}\|=\|{\bf A}(\mathbf{x}-\mathbf{x}^{*})\|\leq\|{\bf A}\|\rho\leq\delta-\frac{\tau(\rho)}{2}$
and for $i=1\ldots,n$,
@@
Let $\mathbf{x}^{1}={\bf A}^{-1}({\bf{b}}+\epsilon\mathbf{e})$ with $\epsilon\leq\rho/\|{\bf A}^{-1}\mathbf{e}\|$. Then we obtain $\|\mathbf{x}^{1}-\mathbf{x}^{*}\|\leq\rho$ and
$n\delta^{2}=f(\mathbf{x}^{*},\hat{\mathbf{y}})$ $>$ $\max_{\|\mathbf{y}\|_{\infty}\leq\delta,\|\mathbf{y}-\delta\mathbf{e}\|_{\infty}\leq\tau(\rho)}\|\mathbf{A}\mathbf{x}^{1}-{\bf{b}}-\mathbf{y}\|^{2}=\|\epsilon\mathbf{e}-\mathbf{y}(\mathbf{x})\|^{2}$
$=$ $(\epsilon-\delta)^{2}n\geq\min_{\mathbf{x}}\max_{\|\mathbf{y}\|_{\infty}\leq\delta,\|\mathbf{y}-\delta\mathbf{e}\|_{\infty}\leq\tau(\rho)}f(\mathbf{x},\mathbf{y}).$
Hence, $(\mathbf{x}^{*},\hat{\mathbf{y}})$ is not a local minimax point of $f$. Moreover, it is easy to see that $(\mathbf{x}^{*},\hat{\mathbf{y}})$ does not satisfy the first order condition (10), since ${\bf A}^{\top}({\bf A}\mathbf{x}^{*}-{\bf b}-\hat{\mathbf{y}})=-\delta{\bf A}^{\top}\mathbf{e}\neq 0$. On the other hand, if $(\mathbf{x}^{\prime},\mathbf{y}^{\prime})$ satisfies the first order condition (10), then from the nonsingularity of ${\bf A}$, we have $f(\mathbf{x}^{\prime},\mathbf{y}^{\prime})=0$, which implies $(\mathbf{x}^{\prime},\mathbf{y}^{\prime})$ cannot be a local minimax point or a global minimax point, since $00=f(\mathbf{x}^{\prime},\mathbf{y}^{\prime})\not\geq\max_{\mathbf{y}\in\Omega}f(\mathbf{x}^{\prime},\mathbf{y})>$ 0.
Note that for nonconvex-nonconcave min-max optimization, a local minimax point must satisfy the first order condition, but a global minimax point is not necessarily a stationary point (Jiang & Chen, 2023).
In this example we present a min-max problem for robust linear least squares problems which does not have a saddle point, a local saddle point and a local minimax point, while it has first order stationary points and global minimax points. However, the set of first order stationary points and the set of global minimax points do not have a common point.
This example also shows that the study of the value function of the inner maximization problem is important. In § 2, we give a proof of closed form expression (5) for the inner value function. Using the inner value function, we prove the existence of a minimax point of the robust nonlinear least squares problem (2). In § 3, we apply the formula to expose some connections between solutions of (2) and those of (1), and give error bounds from any solution of nonlinear least squares problem (1) to the set of minimizers of $\varphi.$ In § 4, we propose a smoothing method with the formula to obtain a complexity bound for finding an $\epsilon$ minimax critical point of a global minimax point of (2). In § 5, we compare least squares solutions and robust least squares solutions in the worst-case, and show the difference of their worst-case residual errors increases linearly as the tolerance of the noise increases. Moreover, we present numerical results of integral equations with uncertain data to show the robustness of solutions of our approach.